Hi would you mind helping me with this question?
My niece Michelle is 6 years old. I am 40. In how many years will i be 3 times as old as her? Let x = the number of years..

Answer by VFBundy(438)

40 + x = 3(6 + x)

40 + x = 18 + 3x

22 = 2x

x = 11